if the diameter of a circle is 3meters, what is the perimeter of the smallest
square that can fully contain the circle?

Answer by oberobic(2304)

If you envision the square touching the circle at 4 points, with the circle inside the square, you can see that the dimensions of the square will be equal to the diameter of the circle. So, it will 3 meters on each side.
The perimeter will be 4*3 = 12 m.
.
Since the perimeter of the square has to be greater than the circumference of the circle, you can check that.
circumference = pi*d
= 3*pi
= 3*3.14, which is < 12.
.
Done.